Present Value of Bonds Payable; Discount Pinder Co. produces and sells high-quality video equipment. To finance its operations, Pinder Co. issued $25,000,000 of five-year, 7% bonds, with interest payable semiannually, at a market (effective) interest rate of 9%. Determine the present value of the bonds payable, using the present value tables in Exhibit 8 and Exhibit 10. Round to the nearest dollar. $
Answer #1
Concepts and reason

Cash flow: Cash flow may be defined as the movement of cash. It comprises of cash inflow and cash outflow. It is used for operating activities of the company. It determines the liquidity status of the company.

Bond: Bond is a long-term financial instrument issued by a company for which the company pays interest to the bond holders and repays the borrowed amount after a specific period of time.

Fundamentals

NPV (Net Present Value): NPV is the net earnings expected by an investor after adjusting “present value” factor from a project. This is the value derived by deducting “present value” cash inflow minus outflow.

Present Value: The value of today’s amount to be paid or received in the future at a compound interest rate is called as present value.

Present Value of Interest factor Annuity (PVIFA): It helps to determine the value in today’s amounts of a series of cash flow to occur at a specified future date.

Present value interest factor (PVIF): The present value of a single sum is the value in today’s amounts of a cash flow to occur at a specified future date. Present value can be determined by discounting the future cash flows.

Calculate the semi-annual coupon rate interest:

Semiannualcouponinterest}=(Facevalueofbond×Couponrate)=$25,000,000×3.5%=$875,000\begin{array}{c}\\\left. \begin{array}{l}\\{\rm{Semiannual coupon }}\\\\{\rm{interest}}\\\end{array} \right\}{\rm{ = }}\left( \begin{array}{l}\\{\rm{Face value of bond }}\\\\ \times {\rm{ Coupon rate}}\\\end{array} \right)\\\\{\rm{ = \$ 25,000,000 }} \times {\rm{ 3}}{\rm{.5\% }}\\\\{\rm{ = \$ 875,000}}\\\end{array}

Therefore the semi-annual coupon rate interest is $875,000.

Calculate the present value of the bonds payable for “P” corporation:

Semi
annual
years
1
Present value
PVF @ PV of Cash
Cash flow formula @
4.5%
flows
4.5%
$875.000 1/(1+0.045)1 0.9569401 $837,3

Therefore the present value of the bonds payable for “P” corporation is $23,021,889.

Ans:

The present value of the bonds payable for “P” corporation is $23,021,889.